OpenStudy (goldeverything):

Somebody please help FAN AND MEDAL WILL BE RECIEVED Which of the following could be the second step in finding the sum negative twenty-two plus negative thirteen using a number line with intervals at every integer? Move twenty-two intervals to the left from zero. Move thirteenintervals to the left from negative twenty-two. Move twenty-two intervals to the right from negative thirteen. Move thirteen intervals to the right from negative twenty-two.

OpenStudy (anonymous):

i think is D)

OpenStudy (goldeverything):

are you sure

OpenStudy (goldeverything):

ok i will go with your gut

OpenStudy (texaschic101):

-22 + (-13) = -22 - 13 = - 35. I believe the answer is B. You would be starting at -22 and moving 13 intervals to the left and ending up at - 35.

OpenStudy (goldeverything):

thank you so much @texaschic101

OpenStudy (texaschic101):

your very welcome :)